# Gravitational Wave emission in Binary Neutron Star early post-merger within a dark environment

### v1

@@ II.1 Binary Neutron Star merger as a damped oscillator
and quite precisely measured in BNS mergers.
We now proceed to generalize the previous idealized oscillator by considering damping may be present due to dissipation. This is physically motivated by the prediction of finite values of NS matter microscopic viscosities for numerous EoS. This has been thoroughly studied in the literature [42, 43]. Models for DM may also include a viscous character by construction, as discussed in Section I.
For the sake of concreteness we will consider that for the strongly interacting NS matter the dissipative force in the Lagrangian dynamics can be generically described as a radial force proportional to $n$-power of radial velocity $v_{r}^{n}=\dot{r}^{n}$ with a density-dependent coefficient, $\alpha_{n}$, proportional to the linear density of ejected matter.
$F_{\rm b,d}$ $=-f_{\rm ejec}\frac{M_{T}}{R_{d}}(\alpha_{1}\dot{r}+\alpha_{2}\dot{r}^{2}),$ (3)
where $\alpha_{1},\alpha_{2}$ are the effective dissipative coefficients of baryonic matter. Similarly, for the dark sector we will consider the possibility of so-called Coulomb damping and viscous damping, both $v_{r}$-dependent with coefficients $\beta_{n}$
$F_{\rm\chi,d}$ $=-\frac{m_{\chi}}{R_{d}}(\beta_{1}\dot{r}+\beta_{2}\dot{r}^{2})=-f_{\chi}\frac{M_{T}}{R_{d}}(\beta_{1}\dot{r}+\beta_{2}\dot{r}^{2}).$ (4)
@@
setting $\alpha_{1}\equiv 0$, or expressed in terms of mass fractions as
@@
We can now apply the Euler-Lagrange formalism to obtain the equations of motion for spatial coordinates $(r,\theta)$ under the form
[…]
with $d_{t}\equiv\frac{d}{dt}$ the time derivative operator. Note that for the dissipative terms, $F_{d}$, the only non-vanishing contribution results from the radial velocity dependence.
$\partial_{\dot{r}}F_{d}=-(2f_{\rm ejec}\alpha_{2}\dot{r}+\beta_{1}f_{\chi}+2f_{\chi}\beta_{2}\dot{r})\frac{M_{T}}{R_{d}}.$ (9)
Finally the equations of motion read
@@ II.1.1 GW Polarizations
Note that the ambient DM present during the merger is characterized by a non-vanishing fraction $f_{\chi}\neq 0$, although we will consider cases where it has either viscous or inviscid nature. This is the minimal assumption that we have introduced as we aim to probe its potential implications on the polarization shape of the waves and spectral properties. More in-depth DM characterization would involve a non-trivial mixing among the two sectors that we will leave for a future work.
To elucidate these effects, we will proceed as follows. First, we will calibrate the DM-free model by obtaining a parameter set adjusting waveforms in the early BNS post-merger from existing numerical relativity simulations. Later on, we will allow a non-zero set of $f_{\chi},\beta_{1,2}$ in order to determine the impact on the results. We will be considering three pivotal scenarios. One where DM exhibits no viscosity ($f_{\chi}\neq 0,\beta_{1}=\beta_{2}=0$), a second one where it showcases viscosity, albeit lower by a factor $\sim 1/10$ than for baryonic matter sourced from stars ($f_{\chi}\neq 0,\beta_{1,2}<\alpha_{2}$), and lastly, a scenario where viscosity is similar or larger as compared to baryonic matter ($f_{\chi}\neq 0,\beta_{2}\gtrsim\alpha_{2}$). To accomplish these tasks, we will employ fitting methodologies grounded in Monte Carlo methods implemented in Python language, see below.
II.1.2 Fitting data sets
@@
Specifically, we extracted simulation data from THC005, THC0032, and THC0040, employing the set of BHBlp [34], SFHo [35], and DD2 [36] EoS, respectively, incorporating nuclear viscosity. The selected dataset shows some spread regarding initial BNS configurations. For example, THC0005 Fig.(1) top panel, relates to a BNS simulation with progenitor masses $m_{1}=1.2M_{\odot}$ and $m_{2}=1.4M_{\odot}$. Stellar structure equations associate for the BHBlp EoS their approximate radii as $R_{1,2}\approx 13.0$ km [34]. THC0032 in Fig.(1) middle panel accurately reproduces the GW signature generated by two NSs with masses $m_{1}=1.20M_{\odot},m_{2}=1.40M_{\odot}$, corresponding to approximate radii of $R_{1}=R_{2}=11.50$ km. Similarly, for THC0040 Fig.(1) lower panel, NS masses are $M_{1}=1.30M_{\odot},M_{2}=1.43M_{\odot}$ with radii $R_{1}=R_{2}=13.0$ km. In our setup we fix somewhat arbitrarily an approximate disk radius $R_{d}\simeq 15.44$ km.
@@
In the model presented, the parameter space encompasses the 10-tuple $(k,\alpha_{2},\omega_{0},r_{0},l,\tilde{f},m_{1},m_{2},f_{\chi},\beta_{n})$ selecting $n=1,2$ or $\beta_{n}\equiv 0$ and $\omega(t_{\rm merger})\equiv\omega_{0}$, $r(0)\equiv r_{0}$. Each parameter holds a distinct physical significance. Specifically, $k$ is the effective spring constant of the elastic force in the oscillator-like gravitational potential in the post-merger. $\alpha_{2},\beta_{n}$ represent the linear density dependence associated with quadratic (or n-power) dissipative forces in Eq. (5) for the ordinary (dark) sector. $\omega_{0}$ is the initial angular velocity of the system, intricately linked to that of the binary system at the moment of merger. $r_{0}$ stands for the initial separation distance between the cores, while the parameter $l$ represents the length where the elastic force between the cores reaches its minimum value. The parameter $\tilde{f}$ stems from $f_{\text{ejec}}\equiv\tilde{f}\frac{M_{T}}{\overline{R}}$ and ${\overline{R}}\equiv\frac{1}{2}\sum_{i=1,2}R_{i}$. This redefinition of the ejected baryonic mass parameter partly accounts for the compactness of the mass-radius relation of the merging NSs and the associated EoS (supposed to be the same for both NSs).
II.1.3 Setup
Using the CoRe database selected cases and the parameter space described in the section II.1.2, we proceed to fit the proposed DO model. To this end, we employ the open-source emcee module in Python [46] which implements the Metropolis-Hastings method as a variant of the Markov Chain Monte Carlo algorithm [46]. Using MCMC we explore the parameter space spanned by the 10-tuple of physical quantities associated with the model, and minimize the distance between CoRe simulation provided main mode $h^{22}$ and the output of the equations of motion in our model shown in Eq.(II.1) from which the solutions $r(t),\theta(t)$ obtained with a Runge-Kutta method are used to construct the GW polarization modes $h_{+}(t)$ and $h_{\times}(t)$ in the form Eq.(14) and Eq.(15).
The simulation data consist of the early post-merger stage of the GW normalized in the temporal coordinate. This normalization significantly facilitates the search for parameter space; however, it must be taken into account when attempting to recover the values of the frequencies associated with this stage. So setting $t_{\rm merger}=0$ the time intervals for each of the data sets THC005, THC0032, and THC0040 will be $\Delta t_{\rm THC005}=3.55$ ms, $\Delta t_{\rm THC0032}=2.36$ ms, $\Delta t_{\rm THC0040}=3.12$ ms, respectively. In addition, we place physically informed priors on some parameters from existing phenomenological studies.
@@ II.1.3 Setup
$m_{1}$ $[M_{\odot}]$ 1.44 1.41 $\pm$ 0.03 1.18 1.18 $\pm$ 0.03 1.48 1.43 $\pm$ 0.11
$m_{2}$ $[M_{\odot}]$ 1.18 1.21 $\pm$ 0.03 1.42 1.40 $\pm$ 0.03 1.26 1.31$\pm$ 0.08
Table 1: The values of the fitted 10-tuple in DM free parameter space fitting selected early post-merger GW waveforms in Fig.(1). The first column (vector) shows the 10-tuple quantities in parameter space along with units. The Best Fit and $\langle x\rangle\pm\sigma(x)$ columns are obtained using the MCMC method. Details are described in Section II.1.3.
@@
Thus, we first identify a set of 10-tuples that fit the early post-merger window for each THC00XX DM-free simulation. The plot of polarizations with fixed parameters to the vectors of the best fit is shown in Fig. (2). More in detail we specify the best fit (as provided by emcee) and the median value along statistical sigma-deviation $\langle x\rangle\pm\sigma(x)$ for each fitted parameter. These values are summarized in the Table (1).
II.1.4 Effective viscosity
Discussing the viscosity of matter in an extreme environment like this proves to be very challenging due to the assumed hybrid nature, ordinary and dark, of matter under study. For strongly interacting matter it is still poorly understood [47] but it is expected it may impact emitted GW patterns [48] while for DM itself we only have some astrophysical and cosmological hints obtained from phenomenology in large-scale structure, where some works [49] describe that CDM may acquire viscous character from decaying to relativistic particles or scattering ordinary matter. In our setting, we will consider the BNS merging event in a quasi-static phase approximation during which the density in the disk during the time window analyzed remains nearly constant so that we can characterize it by an effective kinetic viscosity value. Most works, in practice, consider an interval with different values of the bulk-viscosity coefficient $\sim[0.2,1]\times 10^{30}$ $\rm g/cms$ as in [50]. In that work, authors find that bulk-viscous effects reduce the radiated energy by $\lesssim 1\%$ in the scenario of small viscosity, and at most by $\sim 15\%$ in the scenario of large viscosity. In our calculations, we find fitting scenarios with large/small effective hadronic viscosity parameters, see Table (2).
At this point, it is important to note that in our model dissipative forces are described by characteristic $\alpha_{2},\beta_{n}$ functions where we have made explicit their dependence in linear density. This implies that if we want to compare with microscopic calculations we must first convert them to the well-known dynamic viscosity coefficient. In general, a kinetic viscosity can be defined as $\alpha=\frac{\mu}{\rho}$ where $\mu$ is the dynamic viscosity and $\rho$ is the average volumetric mass density of the merging NS system. In order to convert quantities we write for our model
@@ II.1.4 Effective viscosity
From this, it follows that baryon effective kinetic viscosity $\alpha_{\rm eff}$ with respect to original viscosity $\alpha_{2}$ will be
@@
[…]
@@
Table 2: The effective kinetic viscosity ($\alpha_{\rm eff}$), density ($\rho$) and bulk viscosity ($\mu$) values for the baryonic matter within the BNS merger disk, presented in the cgs system, corresponding to various datasets used in this work.
III Spectral properties
One interesting property characterizing GW emission in BNS mergers appears to be the quasi-universal nature of their frequency spectrum. More in detail, these frequencies can be determined both during the inspiral phase of two compact objects and following their merger. Additionally, it is worth noting that while the precise positioning of peaks in the frequency spectrum may be influenced by the NS matter EoS, this influence can be readily eliminated to establish relationships between peak frequencies and properties of the progenitor stars that are independent of the EoS [52]. From this, one can conclude that the relationship between the BNS inspiral and post-merger stages would provide an opportunity to investigate the NS structure [41] and BH collapse [53] and environmental effects in which the merger occurs or further distortion of gravitational waveforms showing hints of exotic matter types within the progenitor stars.
One example of the mentioned quasi-universal relations can be found after the careful analysis of GW spectra obtained from simulations [54]. The characteristic frequency, commonly denoted by $f_{\rm max}$, representing the frequency at peak amplitude in the post-merger phase is correlated to the tidal deformability $\Lambda$ of individual stars as shown in [55]. For this the authors use a set of simulations with $q\sim 1$ and $M_{T}=2.7M_{\odot}$. Remarkably in that work they find $\rm log_{10}(\frac{f_{\rm max}}{\rm Hz})=3.69652-0.131743\Lambda^{1/5}$ where $\Lambda\equiv\frac{2}{3}\bar{k_{2}}(\frac{\bar{R}}{\bar{M}})^{5}$ and $\bar{R}$ and $\bar{M}$ the radius and mass of the binary system, respectively. $k_{2}$ is the $l,m=2,0$ tidal Love number. The quantities with a bar are defined as averages, i.e., $\bar{A}\equiv\left(A_{1}+A_{2}\right)/2$. They also find a weaker correlation of $f_{\rm max}$ with compactness $C$.
When unequal NS masses are used the relation adopts a slightly different form [1]
$\log_{10}\left(\frac{f_{\max}}{\mathrm{Hz}}\right)\approx 4.2423-0.1546\Lambda^{1/5}-\log_{10}\left(\frac{2\bar{M}}{M_{\odot}}\right).$ (21)
This equation provides us with a way to relate the characteristic frequency of the GW at peak amplitude with the structure of the NSs involved in the merger. More specifically it involves the universality of $\bar{M}f_{\rm max}$. Recent works point out that the ultimate reason for this could be due to the excitation of asteroseismology modes [56].
Figure 3: Panel represents the spectral density $\sqrt{S_{h}(f)}=2\tilde{h}(f)f^{1/2}$ computed for the full post-merger data (thin red line) and for the early post-merger interval after the fitting procedure (thick colored lines) obtained from variation of DM parameters ($f_{\chi}$, $\beta_{1}$) with $\beta_{2}=0$. $\beta_{1}$ is expressed in $\rm 10^{16}(cm/s)^{2}$ units. We also plot sensitivity curves of the NEMO, CE, ET and aLIGO detectors (black and grey curves). Data sets THC005, THC0032, and THC0040 appear in the left, center, right columns, respectively.The lower panel shows the case of $\beta_{1}=0$. In all cases $d=50$ Mpc.
What is most intriguing is that relationships like Eq.(21) could enable us to infer the shape of the frequency spectrum during post-merger phases, relying solely on the information available during the inspiral (cold matter) phase. This would open up the possibility of comparing the complex spectrum of the measured GW in early post-merger with what can be inferred from the more tractable inspiral phase. Any discrepancies between two inspiral spectra for similar BNS events could provide valuable insights into the conditions of the environment in which the merger occurs, including the potential presence of DM or other interesting astrophysical phenomena.

### v2

@@ II.1 Binary Neutron Star post-merger dynamical oscillations
and quite precisely measured in BNS mergers. In our modellization we have assumed that the angular velocity in the disk $\Omega\simeq\dot{\theta}$. We would like to remark here that, as simulations show, see for example [47], differential rotation appears inside the disk. There and after a dozen of ms outer layers at $\sim 15$ km are rotating slightly below Keplerian velocity, while the core rotates more slowly within factors of order unity. However we expect the addition of a viscous effect due to a net fraction of DM will drive even this approximation more realistic and, more importantly, not invalidate our results.
We now proceed to generalize the previous idealized scenario to include DM-induced dissipation. Depending on the conditions of the neutrino transport in BNS mergers dense matter may exhibit relevant microscopic viscous effects. Viscosities for numerous EoS have been thoroughly studied in the literature [48, 49]. We will discuss on this point later on when data sets we use are presented. Models for DM may also include a viscous character by construction, as discussed in Section I and recently shown in BBH simulations where viscous DM facilitates supermassive black holes enough to drop them within a parsec of each other [39].
For the sake of concreteness we will consider the dissipative force $F_{d}$ in the Lagrangian dynamics can be generically described as a radial force proportional to $n$-power of radial velocity $v_{r}^{n}=\dot{r}^{n}$ with a density-dependent coefficient, proportional to the linear density.
@@
where $\alpha_{2},\beta_{1},\beta_{2}$ are the generic dissipative coefficients. $\beta_{1,2}$ are induced by DM interaction while we choose to set a non-zero $\alpha_{2}$ term for baryon. Effects from neutrino leakage scheme and viscous ordinary matter seem to not largely distort the GW waveforms [40] and in that sense we have verified this from our results, see below. In an alternative expression we obtain
@@
We can now apply the Euler-Lagrange formalism to obtain the equations of motion
@@ II.2 GW Polarizations
Note that the ambient DM present during the merger is characterized by a non-vanishing fraction $f_{\chi}\neq 0$, although we will consider cases where it has either viscous or inviscid nature. This is the minimal assumption that we have introduced as we aim to probe its potential implications on the polarization shape of the waves and spectral properties. More in-depth DM characterization would involve a non-trivial mixing involving force carriers among the two sectors that we will leave for a future work.
To elucidate these effects, we will proceed as follows. First, we will calibrate the DM-free model by obtaining a full model parameter set reasonably adjusting waveforms in the early BNS post-merger from selected numerical relativity simulations. Later on, we will allow a restricted non-zero set of $f_{\chi},\beta_{1,2}$ in order to determine the impact on the results. We will be considering three pivotal scenarios. One where DM exhibits no self-interaction viscosity ($f_{\chi}\neq 0,\beta_{1}=\beta_{2}=0$), a second one where it showcases effective viscosity with baryons, albeit lower by a factor $\sim 1/10$ than the previous, sourced from stars ($f_{\chi}\neq 0,\beta_{1,2}<\alpha_{2}$), and lastly, a scenario where self-interaction viscosity is similar or larger than DM-baryon ($f_{\chi}\neq 0,\beta_{2}\gtrsim\alpha_{2}$). To accomplish these tasks, we will employ fitting methodologies grounded in Monte Carlo methods implemented in Python language, see below.
II.3 Fitting data sets
@@
Specifically, we extracted simulation data from THC0005, THC0032, and THC0040, employing the set of BHBlp [42], SFHo [43], and DD2 [44] EoS, respectively. The selected dataset shows some spread regarding initial BNS configurations, grids and use different neutrino schemes: neutrino leakage (LK) for the former two and a more advanced LK+M0 (hereafter M0) for the latter. A key point is neutrino trapping so that, if it occurs, modified-Urca processes can lead to strong bulk viscous dissipation and to damping [52, 53] of the remnant density oscillations possibly leaving a signature in the post-merger GW signal. More advanced schemes, such as M1, found no significant out-of-thermodynamic equilibrium effects on the post-merger. [54]. Thus M1 simulations do not show evidences of enhanced damping of the radial oscillations of the remnant compared to the M0 runs. This suggests that the impact of bulk viscosity cannot be too large. It is however necessary to perform multi-resolution studies in more depth to elucidate this.
Dataset THC0005 Fig.(1) top panel, relates to a BNS simulation with progenitor masses $m_{1}=1.4M_{\odot}$ and $m_{2}=1.2M_{\odot}$, $q=1.166$. Stellar structure equations associate for the BHBlp EoS their approximate radii as $R_{1,2}\approx 13.0$ km [42].
THC0032 in Fig.(1) middle panel accurately reproduces the GW signature generated by two NSs with masses $m_{1}=1.4M_{\odot}$, $m_{2}=1.2M_{\odot}$, corresponding to approximate radii of $R_{1}=R_{2}=11.50$ km. Similarly, for THC0040 Fig.(1) lower panel, NS masses are $m_{1}=1.432M_{\odot}$, $m_{2}=1.3M_{\odot}$, $q=1.1055$ with radii $R_{1}=R_{2}=13.0$ km. In our setup we fix somewhat arbitrarily an approximate disk radius $R_{d}\simeq 15.44$ km.
@@
In the model presented, the parameter space encompasses the 10-tuple $(k,\alpha_{2},\omega_{0},r_{0},l,\tilde{f},m_{1},m_{2},f_{\chi},\beta_{n})$ selecting $n=1,2$ or $\beta_{n}\equiv 0$ and $\omega(t_{\rm merger})\equiv\omega_{0}$, $r(0)\equiv r_{0}$. Each parameter holds a distinct physical significance. Specifically, $k$ is the effective spring constant of the elastic force defined for the gravitational potential in the post-merger. $\alpha_{2},\beta_{n}$ are associated with dissipative forces in Eq. (3) for the ordinary (dark) sector. $\omega_{0}$ is the initial angular velocity of the system, intricately linked to that of the binary system at the moment of merger. $r_{0}$ stands for the initial separation distance between the cores, while the parameter $l$ represents the length where the elastic force between the cores reaches its minimum value. The parameter $\tilde{f}$ stems from $f_{\text{ejec}}\equiv\tilde{f}\frac{M_{T}}{\overline{R}}$ and ${\overline{R}}\equiv\frac{1}{2}\sum_{i=1,2}R_{i}$. This redefinition of the ejected baryonic mass parameter partly accounts for the compactness of the mass-radius relation of the merging NSs and the associated EoS (supposed to be the same for both NSs).
II.4 Setup
Using the CoRe database selected cases and the parameter space described in the section II.3, we proceed to fit the proposed DO model. To this end, we employ the open-source emcee module in Python [55] which implements the Metropolis-Hastings method as a variant of the Markov Chain Monte Carlo algorithm [55]. Using MCMC we explore the parameter space spanned by the 10-tuple of physical quantities associated with the model, and minimize the distance between CoRe simulation provided main mode $h^{22}$ and the output of the equations of motion in our model shown in Eq.(II.1) from which the solutions $r(t),\theta(t)$ obtained with a Runge-Kutta method are used to construct the GW polarization modes $h_{+}(t)$ and $h_{\times}(t)$ in the form Eq.(9) and Eq.(10).
The simulation data consist of the early post-merger stage of the GW normalized in the temporal coordinate. This normalization significantly facilitates the search for parameter space; however, it must be taken into account when attempting to recover the values of the frequencies associated with this stage. So setting $t_{\rm merger}=0$ the time intervals for each of the data sets THC0005, THC0032, and THC0040 will be $\Delta t_{\rm THC0005}=3.55$ ms, $\Delta t_{\rm THC0032}=2.36$ ms, $\Delta t_{\rm THC0040}=3.12$ ms, respectively. In addition, we place physically informed priors on some parameters from existing phenomenological studies.
@@ II.4 Setup
$m_{1}$ $[M_{\odot}]$ 1.44 1.41 $\pm$ 0.03 1.42 1.42 $\pm$ 0.03 1.48 1.43 $\pm$ 0.11
$m_{2}$ $[M_{\odot}]$ 1.18 1.21 $\pm$ 0.03 1.18 1.18 $\pm$ 0.03 1.26 1.31$\pm$ 0.08
Table 1: The values of the fitted 8-tuple fitting selected early post-merger GW waveforms in Fig.(1) with no DM $f_{\chi}=\beta_{1,2}\equiv 0$. The first column (vector) shows the 10-tuple quantities in parameter space along with units. The Best Fit and $\langle x\rangle\pm\sigma(x)$ columns are obtained using the MCMC method. Details are described in Section II.4.
@@
Thus, we first identify a set of 10-tuples that fit the early post-merger window for each THC00XX simulation assuming no DM. We nevertheless allow at this point a residual non-zero viscous force $\propto\alpha_{2}$ related to the baryon content. The plot of polarizations with fixed parameters to the vectors of the best fit is shown in Fig. (2). More in detail we specify the best fit (as provided by emcee) and the median value along statistical sigma-deviation $\langle x\rangle\pm\sigma(x)$ for each fitted parameter. These values are summarized in the Table (1).
Elucidating the microscopic effect of the viscosity of matter in an extreme environment like this proves to be very challenging due to the assumed hybrid nature, ordinary and dark, of matter under study. We will assume that the self-interacting DM polluting the standard NS matter may effectively impact emitted GW patterns while for self-interacting DM sector we only have some astrophysical and cosmological hints obtained from phenomenology in large-scale structure where some works [56] describe that CDM may acquire viscous character from decaying to relativistic particles or scattering ordinary matter. Recent simulations may also indicate additional effects [39]. In our setting, we will consider the BNS merging event in a quasi-static phase approximation during which the density in the disk during the time window analyzed remains nearly constant so that we can characterize it by an effective hybrid kinetic viscosity value.
At this point, it is important to note that in our model dissipative forces are described by characteristic $\alpha_{2},\beta_{n}$ functions where we have made explicit their dependence in linear density.
In general, a kinetic viscosity can be defined as $\alpha=\frac{\mu}{\rho}$ where $\mu$ is the dynamic viscosity and $\rho$ is the average volumetric mass density of the merging NS system. In order to convert quantities we write for our model
@@
From this, it follows that
@@
Given that we have taken an approximate radius $R_{d}\simeq 15.44$ km, we can place bounds on the height of the matter disk as $2R_{S}<h<2R_{d}$ such that it is smaller than disk diameter and greater than twice the Schwarzschild radius, $R_{S}$, of the associated mass of the nuclei oscillating inside, we fix $h\sim 15$ km. Thus, we take the characteristic distance of the system such that it is on the order of twice the Schwarzschild radius, which holds for all the cases studied in this article as $l_{c}\lesssim 2G\bar{M}/c^{2}\sim 9$ km, $\bar{M}\equiv\left(m_{1}+m_{2}\right)/2$. The conversion can be expressed as $\mu=\alpha_{\rm eff}\rho$. Using the values collected in Table 1 and based on the above assumptions, we see the DM-baryon interaction is rather weak not only for data sets THC0005, THC0032 but also for THC0040 where M0 neutrino scheme is used. In any case our setting must be refined to make any quantitative claim on the viscous effects from hadronic matter.
@@
Table 2: The effective kinetic viscosity ($\alpha_{\rm eff}$), density ($\rho$) and bulk viscosity ($\mu$) values for hybrid matter within the BNS merger disk, presented in the cgs system, corresponding to various datasets used in this work.
III Dark Matter imprints on the spectral distribution
We now assess further qualitative changes from the GW patterns as obtained within our model for nearly symmetric BNS mergers. We do not attempt, as mentioned, precisely characterizing the impact of DM but rather spotting how viscous DM may influence the spectral distribution.
On more technical grounds, one interesting property characterizing GW emission in BNS mergers appears to be the quasi-universal nature of their frequency spectrum for systems with $q=1$. However, some authors [37] claim in asymmetric systems tidal disruption of the NS cores ruins this picture. In their Fig. 11 and Fig. 12 in that work they show that the $f_{1}$ peak is less clearly detectable as BNS mass asymmetry $q$ grows. Their asymmetric cases departing from ($\eta=0.2500)$ or equivalently $q=1$ show the peak smears off. In our analysis we use $\eta=0.2494,0.2485$ so we expect our qualitative conclusions still hold.
Figure 3: Panel represents the spectral density $\sqrt{S_{h}(f)}=2\tilde{h}(f)f^{1/2}$ computed for the full post-merger data (thin red line) and for the early post-merger interval after the fitting procedure (thick colored lines) obtained from variation of DM parameters ($f_{\chi}$, $\beta_{1}$) with $\beta_{2}=0$. $\beta_{1}$ is expressed in $\rm 10^{16}(cm/s)^{2}$ units. We also plot sensitivity curves of the NEMO, CE, ET and aLIGO detectors (black and grey curves). Data sets THC0005, THC0032, and THC0040 appear in the left, center, right columns, respectively. The lower panel shows the case of $\beta_{1}=0$. In all cases $d=50$ Mpc.